3. The Pew Research Center wishes to estimate the proportion of adults who are aware of Instagram. How many adults must the marketing firm randomly survey if they want to be 96% confident that the proportion of adults who are aware of Instagram is within 3% of the true proportion of adults who are aware of Instagram? Enter answer 4 The mean age of a sample of 40 college students is 23.95 with a standard deviation of 2.55. Assume that prior studies indicate that the standard deviation of college students ages is 2.04 Construct a 99% confidence interval estimate of the mean age of all college students. What is the upper bound of the confidence interval? (Round to the nearest thousandth.)
Answer #1

3)

here margin of error E = 0.03
for96% CI crtiical Z          = 2.05
estimated prop.=p= 0.50
reqd. sample size n=      p*(1-p)*(z/E)2= 1168

4)

sample mean 'x̄= 23.950
sample size    n= 40.00
std deviation σ= 2.040
std errror ='σx=σ/√n= 0.3226
for 99 % CI value of z= 2.58
margin of error E=z*std error = 0.831
lower bound=sample mean-E= 23.118
Upper bound=sample mean+E= 24.782